OpenStudy (tester97):

a doctors office schedules 15-minute appointments and half-hour appointments for weekdays. the doctor limits these appointments to, at most, 30 hours per week. write an inequality to represent the number of 15-minute appointments x and the number of half-hour appointments y the doctor may have in a week.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

answer is 15x+30y<=1800 <= equals < with a line under it

OpenStudy (tester97):

oh ok thanks so much

OpenStudy (tester97):

but where does the 1800 come from?

OpenStudy (anonymous):

lol....u want a full explanation, i ca tell u this its 100% right

OpenStudy (anonymous):

????

OpenStudy (tester97):

no i was just wondering thats all :P

OpenStudy (tester97):

so where does it come from? do u multiply 30 times 60?

OpenStudy (anonymous):

Yes, he multiplied 30*60 (as it is 60min in an hour). You could convert 15min and 30min into hours, leaving 30 as is. \[\frac{1}{4}x+\frac{1}{2}y \leq 30\] By multiplying every term with 60( to convert from hours to minutes) we get \[15x+30y \leq 1800\]